OpenStudy (anonymous):

How can u keep simplifying an expression if they have nothing left in common. I'm asked to simplify 5n^2+21n-4

OpenStudy (shubhamsrg):

Do you know a quadratic eqn like px^2 + qx + r =0 can be expressed in the form px^2 - (a+b)x + ab = 0 ? Where, a and b are the roots of the eqn, i.e. this quadratic eqn will also represent (x-a)(x-b) =0

OpenStudy (shubhamsrg):

I might not be very clear, please tell me if that is the case.

OpenStudy (solomonzelman):

you can't simplify this expression. \(\Large\color{orangered}{ \bf (21)^2-4(5)(-4) }\) \(\Large\color{orangered}{ \bf 441-4(5)(-4) }\) \(\Large\color{orangered}{ \bf 441+80 }\) \(\Large\color{orangered}{ \bf 521 }\) \(\Large\color{red}{ \bf 484=(±22)^2 }\) \(\Large\color{red}{ \bf 529=(±23)^2 }\) the discriminant is not a perfect square so you can't (simply) factor this either.

OpenStudy (shubhamsrg):

Oh right, I didn't notice. Can you please check the question ? Is it +4 ?

OpenStudy (solomonzelman):

yeah, with +4 it would be easier :)

OpenStudy (anonymous):

how do u type in red

OpenStudy (solomonzelman):

I will message you.

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

Try to put something up

OpenStudy (solomonzelman):

\(\Large\color{purple}{ \bf H }\)\(\Large\color{blue}{ \bf e }\)\(\Large\color{royalblue}{ \bf l }\)\(\Large\color{green}{ \bf l }\)\(\Large\color{red}{ \bf o }\)\(\Large\color{orangered}{ \bf ! }\)

OpenStudy (anonymous):

\(\Large\color{red}{ hello }\)

OpenStudy (anonymous):

\(\Huge\color{red}{ HELLO }\)

OpenStudy (solomonzelman):

yes, you can use a couple fonts and sizes in front of the phrase in LaTeX. Sizes \tiny \scriptsize \small \normalsize \large \Large \LARGE \huge \Huge \(\Large\color{orangered}{ \bf ! }\) fonts \tt \it \bf \rm \cal \scr \frak

OpenStudy (anonymous):

Thank you so much !

OpenStudy (solomonzelman):

Anytime